OpenStudy (anonymous):

Factor the polynomial 12x^2 - 4x

OpenStudy (anonymous):

Can you see what is common between the terms 12x^2 and 4x?

OpenStudy (anonymous):

GCF: 2?

OpenStudy (anonymous):

okay 2 is one but there is a greater common factor

OpenStudy (anonymous):

4

OpenStudy (anonymous):

perfect so 4 is common between both terms. now there is also an x that is common So we want to take a 4x out so 12x^2-4x=4x( - ) If I take the 4x out what will I be left with?

OpenStudy (anonymous):

12x - 4?

OpenStudy (anonymous):

Okay that's close you have taken the x term out correctly but we also need to take out a factor of 4. What will we get?

OpenStudy (anonymous):

We'd get 8x?

OpenStudy (anonymous):

Okay so right now we have 12x^2-4x=x(12x-4) to take out the factor of 4 i need to divide what is inside the brackets by 4 So 4x(12x/4-4/4)=4x(3x-1)

OpenStudy (anonymous):

Ohhhh! I see, I get it!

OpenStudy (anonymous):

My mistake

OpenStudy (anonymous):

No worries, would you like a question to try out?

OpenStudy (anonymous):

Yes, I have another question on my worksheet.

OpenStudy (anonymous):

18x^2 - 10x

OpenStudy (anonymous):

Okay would you like to try this first and I'll check it :)

OpenStudy (anonymous):

okay. one sec.

OpenStudy (anonymous):

So, 18x^2-10x = 2x(18x-10) 2x(9x-5) ?

OpenStudy (anonymous):

Perfect, nice work!

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

Can you please help me with one more problem?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

10x^3-12x^2+8x

OpenStudy (anonymous):

You want to use the exact same method but now you have three terms. So what factors are you going to take out?

OpenStudy (anonymous):

10x^3-12x^2+8x = 2x(10x-12+8) = 2x(5x-6+4) ?

OpenStudy (anonymous):

very close all the number are right 10x^3-12x^2+8x=2x(5x^3/x-6x^2/x+4x/x) ==2x(5x^2-6x+4)

OpenStudy (anonymous):

I have to head off to work, but I'll check this afterwards. Good luck :D

OpenStudy (anonymous):

Thank you so much for everything :)